# Versatile Motion-Language Models for Multi-turn Interactive Agents

### v2

@@ 4.2 Architecture
Our architecture for modeling and generating interactive motions consists of three primary components: encoders (tokenizers), a large language model block, and decoders. This design allows for the integration of both motion and text data within a unified framework. For motion, we use a residual vector quantized variational auto-encoder (RQ-VAE (Lee et al., 2022; Guo et al., 2024a)) as a tokenizer. Vector quantized variational auto-encoders (Van Den Oord et al., 2017) are effective, but their quantization causes information loss and reduces reconstruction quality, which is critical for accurately modeling interactive motions. The motion encoder $\mathcal{E}_{M}$ applies 2D convolutions to motion features along the time axis, converting motion pairs ${\mathbf{m}_{a},\mathbf{m}_{b}}$ into latent vectors $\{\mathbf{z}^{1:L}_{a},\mathbf{z}^{1:L}_{b}\}=\mathcal{E}_{M}(\{\mathbf{m}^{1:M}_{a},\mathbf{m}^{1:M}_{b}\})$, where $M$ is a motion length and $L=M/l$ with down-sample rate $l$. Then the latent vectors $\mathbf{z}$ are quantized by RQ-VAE as an ordered $D$ discrete codes:
where $\mathcal{C}$ is the codebook, $K=|C|$, and $k^{i}_{d}$ is code of $\mathbf{z}$ at timestep $i$ and depth $d$. These discrete codes form a motion vocabulary. For text, we use a standard text tokenizer to process textual instructions and descriptions into tokens.
Tokens are then proceeded to the language model block, which serves as the central processing unit. In this work, we have utilized the LLaMA-3.1-8B (Dubey et al., 2024) model as a base model. We integrate motion tokens with text tokens using a unified vocabulary, which combines both the text and motion vocabularies into one, with special tokens added to mark the start and end of the motion sequences. This shared token space enables the model to efficiently process and generate both modalities for motion-related tasks. Interactive motion is represented as $X_{m}=\{k^{1;a}_{1:D},k^{1;b}_{1:D},\cdots,k^{L;a}_{1:D},k^{L;b}_{1:D}\}$, where $X_{m}$ is a sequence of motion represented in unified vocabulary and $k^{i;a}_{1:D}\in[K]^{D}$ is the $i$-th token of motion $a$.
Finally, the decoding stage reverses the encoding process. For motion, the decoder $\mathcal{D}_{M}$ projects the quantized features $\hat{\mathbf{z}}^{i}=\sum_{d=1}^{D}\mathbf{e}(k_{d}^{i})$, where $\mathbf{e}$ is codebook embedding, back into motion sequences using 1D convolution. Text decoding follows standard language model decoding procedures.
@@ 4.3 Training
Figure 3: Method Overview. Stage 1 involves training a motion tokenizer that encodes and decodes interactive motion data. In Stage 2, we pre-train the model by integrating motion and text data, allowing it to learn the alignment between text and motion. Stage 3 focuses on Instruction Tuning, fine-tuning the model to follow instructions and improve its responsiveness to conversational cues.
@@
The first stage is to train a motion tokenizer composed of an encoder, decoder, and quantizer. We followed the original objective functions from Lee et al. (2022) to train this model, minimizing the reconstruction loss, the codebook loss to align the encoder’s outputs with the codebook, and the commitment loss to ensure encoder consistency. Once the encoder and decoder are optimized, we maintain this model freezed during the rest of the training stage.
Pre-training Strategy
In the pre-training stage, we train a pre-trained large language model to align motion representations with textual representations. We design tasks including motion-to-text, text-to-motion, motion prediction, and reaction generation to train the model, leveraging paired datasets like Inter-X (Xu et al., 2024a) and InterHuman (Liang et al., 2024). Using the template from Jiang et al. (2023), we create input sequences $y$ from motion sequences $X_{m}$ and the corresponding motion caption. Since both motion tokens and text tokens are discrete, we train our model with the general language modeling next-token prediction objective: $\mathcal{L}=-\log\sum^{T}p_{\theta}(y_{i}|y_{<i})$, where $T$ is the length of the multi-modal sequence and $i$ only counts when the text token appears at position $i$. To improve training efficiency, we train the LLM using a low-rank adaptor (LoRA) (Hu et al., 2022), similar to Ge et al. (2024). We then merged the LoRA parameters to the LLM backbone for further training. Furthermore, due to a limited number of interactive motion data, we also leverage larger single motion-text datasets from Motion-X (Lin et al., 2024). This offers prior knowledge of how individual motions are described in language, enhancing the model’s ability to align motions with corresponding textual descriptions.
Instruction-tunning with Inter-$\text{MT}^{2}$ Data
In this stage, we aim to enhance the model’s ability to follow a wide range of instructions presented in a conversational format. We utilize Inter-$\text{MT}^{2}$ dataset combined with single-turn data from prior interactive motion datasets (Xu et al., 2024a; Liang et al., 2024). We follow instruction templates from MotionGPT (Jiang et al., 2023), to format the input and output for single-turn data. The multi-modal sequence $y$ consists of user instructions and corresponding responses, formatted as $y=(X^{1}_{u},X^{1}_{a},X^{2}_{u},X^{2}_{a},\cdots)$. The training objective remains the same as in the pre-training stage, with user instructions omitted in the loss function.
4.4 Advanced Downstream Interactive Motion Tasks
After training, our model can perform complex reasoning and generate interactive motions within multi-turn conversations. To verify this, we introduce two additional tasks requiring advanced capabilities: motion reasoning and editing. Motion reasoning involves predicting past or future events, or reasoning about current motions, based on prior conversational data. This task requires the model to understand the context of the conversation, interpret how the given motion fits within that context, and adjust its reasoning accordingly. In the motion editing task, we focus on altering a person’s persona or shifting scenarios, such as emotions or relationship dynamics. This adds complexity, as changes to one person’s behavior affect the other’s motion. The model must edit the target motion while maintaining contextual coherence, requiring a deep understanding of social dynamics.
@@ 5 Experiments
In our experiments, we evaluated VIM’s ability to generate detailed motion-based chat responses, requiring complex reasoning about interactive motions, alongside traditional motion-related tasks. We focused on two main questions: first, whether the model can reason effectively about interactive motions, such as refining motions in editing tasks or generating contextually accurate narratives in motion reasoning. Second, we evaluated whether the training with Inter-$\text{MT}^{2}$ dataset improves the model’s performance in text-to-motion, motion-to-text, and reaction generation tasks.
@@ 5.1 Evaluation Tasks and Baselines
Motion reasoning involves predicting past or future events or interpreting current motions using prior conversational context. We utilize powerful LLMs, i.e., GPT-4o (OpenAI, 2024) to assess the content alignment, naturalness, and logical coherence of the generated textual responses. Content alignment evaluates how accurately the text reflects the given motion data, logical coherence checks the consistency and reasoning accuracy of inferences made about past or future events, and naturalness evaluates the fluency of generated texts, with rating each metric on a 10-point scale. In addition, we utilized linguistic metrics like Rouge-L (Lin (2004)), METEOR (Banerjee & Lavie (2005)), and MAUVE (Pillutla et al. (2021)), to quantitatively assess the relevance, accuracy, and naturalness of the generated responses compared with labeled texts in Inter-$\text{MT}^{2}$ test dataset. We present the results on motion reasoning in §5.2.
@@
The goal of motion editing is to modify a reference motion based on user instructions. We conducted within-subject user studies to compare edited motion samples, with participants rating them on three metrics: content similarity, instruction alignment, and motion quality, using a 5-point Likert scale, following Goel et al. (2024). Content similarity evaluates whether the edited motion preserves the original meaning of the source motion, while instruction alignment assesses how accurately the edited motion follows the given command. The study involved 30 participants, each evaluating five samples from a set of 30 randomly selected test samples. Participants evaluated four baselines and our method, viewing randomly shuffled motion outputs side by side and providing feedback on all metrics. We also employed data-driven metrics, such as Frechet Inception Distance (FID) and mean per joint position error (MPJPE) in meters, to evaluate the quality of the generated edited motion against the labeled motions in the Inter-$\text{MT}^{2}$ test set, following Goel et al. (2024). The results on motion editing is shown in §5.3.
@@
For standard motion-related tasks, we evaluated the proposed method in three traditional motion-relevant tasks in interactive motions: motion-to-text, text-to-motion, and reaction generation, in the union of the test set in InterHuman (Liang et al., 2024) and Inter-X (Xu et al., 2024a) datasets. To evaluate the text-motion matching score, we report the retrieval precision based on the feature space of retrieval models (Petrovich et al. (2023)). This evaluates the accuracy of matching between texts and motions using Top 3 retrieval accuracy with a fixed batch of 32. The quality of motion was measured by Frechet Inception Distance (FID), which measures the distance of feature distribution between motion data and generated motion. In addition, we measure the mean per joint position error (MPJPE) in meters to evaluate the accuracy of the reaction motion. The results on motion-related tasks are shown in §5.4.
@@
Since our interactive multi-turn scenarios and tasks are novel, there is no exact comparison method. We consider and compare reasonable baselines that handle both motion and texts as input and output. We first employ a two-stage approach using off-the-shelf methods. For the motion reasoning task, we convert all motions into text descriptions using the state-of-the-art motion-to-text method TM2T (Guo et al., 2022). These textual descriptions are then used for text-based reasoning with large language models such as GPT-4o (OpenAI, 2024) and LLaMA-3.1-8B (Dubey et al., 2024). For the motion editing task, we combine the text converted by TM2T with the editing text command, and generate the edited motion using the off-the-shelf text-to-motion method InterGen (Liang et al., 2024). As baselines for the unified approach, we leverage MotionGPT (Jiang et al., 2023) framework with the following configurations: (1) $\text{MotionGPT}^{*}$: a modified MotionGPT fine-tuned on interaction data with instruction templates; (2) $\text{MotionGPT}^{*}_{I}$: $\text{MotionGPT}^{*}$ enhanced with Inter-$\text{MT}^{2}$ dataset; (3) VIM w/o Inter-$\text{MT}^{2}$: our method fine-tuned with instruction templates from MotionGPT, but without Inter-$\text{MT}^{2}$ data.
@@ 5.2 Motion Reasoning
Methods LLM-Assisted Linguistic Metrics
Coh. $\uparrow$ Align. $\uparrow$ Nat. $\uparrow$ ROUGE-L METEOR MAUVE
two-stage approach
TM2T + LLaMA-3.1-8B 3.852 3.050 6.348 0.158 0.226 0.009
TM2T + GPT-4o 4.266 3.455 6.790 0.162 0.227 0.019
unified approach
$\text{MotionGPT}^{*}$ 1.855 1.303 3.574 0.113 0.096 0.005
$\text{MotionGPT}^{*}_{I}$ 3.690 3.160 5.291 0.207 0.218 0.417
VIM w/o Inter-$\text{MT}^{2}$ 2.770 2.141 4.968 0.155 0.145 0.004
VIM (Ours) 5.252 4.511 6.981 0.239 0.260 0.794
@@
In the motion reasoning task, conversations about two interactive motions are examined to assess the model’s ability to deduce past or future events and comprehend the motivations driving the motions. The experimental results in Table 1 demonstrate that our unified model, VIM, significantly outperforms two-stage approaches across all LLM-assisted and linguistic metrics. Specifically, VIM achieves improvements with performance increases exceeding 1.9 points in logical coherence, 1.1 points in content alignment, and nearly 0.2 points in naturalness compared to the best two-stage model. The baseline models trained solely on text-motion pair datasets, such as VIM w/o Inter-$\text{MT}^{2}$ and $\text{MotionGPT}^{*}$, show limited reasoning capabilities. Although $\text{MotionGPT}^{*}_{I}$, which incorporates Inter-$\text{MT}^{2}$ datasets, exhibits improved performance compared to baselines trained without Inter-$\text{MT}^{2}$, it still does not match the effectiveness of the two-stage approaches.
The improved performance of our unified model, VIM, over two-stage approaches, appears to result from two key factors: error accumulation and interpretation ambiguity. First, in two-stage models, errors can accumulate; if the motion captioning model generates incorrect motion captions, those mistakes carry over to the second stage, reducing content alignment and coherence. In contrast, VIM’s unified architecture integrates motion encoding and reasoning in a single framework, minimizing error propagation. Second, interpreting motions is not always straightforward, with multiple ways to understand and describe the same motion. In two-stage models, mapping the motion to a single caption for the second stage can lead to more contextually accurate reasoning of the given scenarios or contexts. Our unified model, however, is built to recognize these varied interpretations and generate reasoning that is more contextually accurate. Figure 4 showcases it’s ability to dynamically adjust interpretations and responses by incorporating context from previous conversations.
@@ 5.3 Motion Editing
Table 2: Data-driven evaluation in motion editing in Inter-$\text{MT}^{2}$ test set.
@@
Figure 5: User subject study results for motion editing.
We aim to validate the hypothesis that people will perceive the generated edited interactive motion from the proposed method to be more content-consistent, instruction-aligned, and better quality, through user subject studies. To analyze the results, we conducted a repeated-measures multivariate analysis of variance on the rated measures. We observed that methods significantly affect the user’s perception of all dimensions; $F(4)=4.591,p=0.002,\eta^{2}=0.137$ for content similarity, $F(4)=7.134,p=0.000,\eta^{2}=0.197$ for instruction alignment, and $F(4)=4.781,p=0.001,\eta^{2}=0.142$ for motion quality, with all $\alpha=0.05$. The estimated marginal mean of the rated score is reported in Figure 5111We plotted the difference in a post hoc pairwise comparison of the proposed method only. We denote * as $0.050.01<p<$ 0.05, ** as $p<0.01$, and *** as $p<0.001$. The error bars represent 95% confidence intervals. . The results show that the proposed method had better instruction alignment, quality, and content consistency across other baselines with significant differences.
During post hoc pairwise comparisons, we identified a significant difference, with our proposed method outperforming the two-stage model (TM2T (Guo et al., 2022) with InterGEN (Liang et al., 2024)) in content similarity ($p=0.017$) and instruction alignment ($p=0.010$). The two-stage model showed lower content similarity due to motion-to-text conversion errors, leading to unintended motions, whereas our unified framework avoids such error accumulation. Additionally, the two-stage model struggled with instruction alignment because InterGEN was trained to generate motions from captions, limiting its ability to adapt to varying personas or contexts. In contrast, our method, trained on diverse instructions and tasks, demonstrated superior reasoning and adaptability, resulting in more accurate motion generation based on instructions and source motions.
@@
In post hoc pairwise comparisons with $\text{MotionGPT}^{*}_{I}$, we observed significant differences, with our proposed method performing better in content similarity ($p=0.005$), instruction alignment ($p<0.0005$), and motion quality ($p=0.009$). This suggests that the VQ-VAE-based tokenizer and conditional generation model negatively impacted performance. Additionally, compared to VIM w/o Inter-$\text{MT}^{2}$, there were significant differences in content similarity ($p=0.010$) and instruction alignment ($p=0.001$), indicating that without Inter-$\text{MT}^{2}$ data, the model struggles to control motion based on context and instructions. We also evaluated the proposed method using data-driven metrics, including FID and MPJPE, as shown in Table 2. The proposed method outperforms the baselines on both measures, which is consistent with the results from user studies. Figure 6 illustrates the generated edited motions based on the source motion and instruction.

### v5

@@ 4.2 Architecture
Our architecture for modeling and generating interactive motions consists of three primary components: motion tokenizer, large language model (LLM), and motion decoder. This design allows for the integration of both motion and text data within a unified framework. The overview of VIM’s architecture is shown in Figure 3.
To enable the LLM to interpret interactive motions, we first tokenize the motion sequences. We utilize RQ-VAE [21] as a tokenizer to reduce the information loss during the quantization, similarly to the approach in [15].
Figure 4: Tokenization of interactive motions.
The motion encoder $\mathcal{E}_{M}$ applies 2D convolutions to motion features along the time axis, converting motion pairs $\{\mathbf{m}_{a},\mathbf{m}_{b}\}$ into latent vectors $\{\mathbf{z}^{1:L}_{a},\mathbf{z}^{1:L}_{b}\}$, $L=M/l$ with down-sample rate $l$. Each latent vector $\mathbf{z}^{i}$ is quantized into an ordered set of D discrete codes, $\mathcal{RQ}(\mathbf{z}^{i};\mathcal{C},D)=(k^{i}_{1},\cdots,k^{i}_{D})\in[K]^{D}$, where $\mathcal{C}$ is the code book with $K=|\mathcal{C}|$, and $k^{i}_{d}$ is the code of $\mathbf{z}$ at timestep $i$ and depth $d$. These tokens, combined with special tokens indicating the start and end of motions, constitute the motion vocabulary. For text inputs, we utilize a standard text tokenizer compatible with the LLM.
Subsequently, the quantized tokens are provided to the LLM block, which serves as the central processing component. In this work, we initialize VIM with the pretrained LLaMA-3.1-8B [9]. The motion vocabulary and text vocabulary of the LLM are integrated into a unified vocabulary, allowing the model to to efficiently process and generate both modalities. Interactive motion is represented as $X_{m}=\{k^{1;a}_{1:D},k^{1;b}_{1:D},\cdots,k^{L;a}_{1:D},k^{L;b}_{1:D}\}$, where $X_{m}$ denotes the motion sequence encoded in the unified vocabulary, and $k^{i;a}_{1:D}\in[K]^{D}$ is the $i$-th token of motion $a$.
Finally, to visualize the generated motion tokens, we use the motion decoder of the RQ-VAE. The decoder projects the quantized features $\hat{\mathbf{z}}^{i}=\sum_{d=1}^{D}\mathbf{e}(k_{d}^{i})$, converting them back into motion sequences.
@@ 4.3 Training
We describe the training strategy in VIM, to convert a large language model into an interactive motion-language model.
@@
The motion tokenizer consists of an encoder, decoder, and quantizer. We followed the original objective functions from [21], minimizing the reconstruction loss, the codebook loss to align the encoder’s outputs with the codebook, and the loss of commitment to ensure the consistency of the encoder. After training the encoder and decoder, we freeze their parameters throughout the rest of the training stage.
Pre-training for Cross-modal Motion-Text Alignment
The goal of this stage is to enable the large language models (LLMs) to process and generate interaction motion tokens effectively. To achieve this, we continuously pre-train LLMs using paired interaction motion-text datasets, such as Inter-X [39] and InterHuman [22], across various tasks including motion-to-text, text-to-motion, motion prediction, and reaction generation.
For each task, we construct sequences $y$ that combine motion sequences with their corresponding captions and train with a next-token prediction objective $\mathcal{L}=-\log\sum^{T}p_{\theta}(y_{i}|y_{<i})$. To improve training efficiency, we employ LoRA adaptor [17], similar to [11], and merge its parameters to the LLM backbone. Furthermore, due to a limited number of interactive motion data, we also leverage a subset of single-person motion-text datasets from Motion-X [23]. This additional single-person data offers prior knowledge of how the individual motions are described in language, enhancing the model’s ability to align motions with textual descriptions.
Instruction-tuning with Inter-$\text{MT}^{2}$ Data
In this stage, we aim to enhance the model to extend beyond understanding and generating single-turn interaction motions and focusing on handling diverse and complex instructions presented through multi-turn conversational scenarios. Similar to the pre-training stage, We adopt a next-token prediction training objective for training. The instruction-tuning sequences are composed of user interactions paired with corresponding responses, integrating tokens from a unified vocabulary that covers texts, motions, or both modalities. We also leverage the Inter-$\text{MT}^{2}$ dataset along with single-turn interaction data from existing motion datasets [39, 22], formatted according to the instruction template of [18].
@@ 5 Experiments
In this section, we evaluate the effectiveness of VIM, particularly focusing on its capability to accurately understand and generate interactive motions in complex, multi-turn conversational scenarios involving both motions and text modalities. To extensively validate our approach, we compare VIM against several specialized baseline methods, each explicitly designed for individual tasks. This allows us to understand the performance and versatility of VIM. Additionally, we investigate the contribution and effectiveness of our proposed dataset, Inter-$\text{MT}^{2}$, showing how it enhances VIM’s ability to process and generate interactive motion and texts. We also provide qualitative video results generated by VIM in the supplementary material.
@@ 5.1 Evaluation Tasks and Baselines
We introduce a motion reasoning task to validate the model’s ability to comprehend interactive motions and text queries. Motion reasoning involves predicting past or future events, or reasoning about current motions, based on prior conversational data. This task requires the model to understand the context of the conversation, interpret how the given interactive motion fits within that context, and adjust its reasoning accordingly. We utilize LLMs-based evaluator, specifically GPT-4o [27], to assess the content alignment, naturalness, and logical coherence of the generated textual responses. Content alignment evaluates how accurately the text reflects the given interactive motions, logical coherence checks the consistency and reasoning accuracy of inferences made about past or future events, and naturalness evaluates the fluency of generated texts, with rating each metric on a 10-point scale. Additionally, we employ linguistic metrics, such as METEOR [4], and MAUVE [30] to quantitatively evaluate relevance and fluency against 2002 labeled samples from the Inter-$\text{MT}^{2}$ test set. We present the results on motion reasoning in §5.2.
@@
In the motion editing task, the model modifies the given motion based on a person’s persona or scenario, e.g., emotions or relationship dynamics, which adds complexity as changes in one individual affect the other’s motion. Unlike single-person motion editing [2, 13], the task that edits interactive motions should consider preserving contextual coherence and social dynamics. We evaluated the methods on 1445 samples from the Inter-$\text{MT}^{2}$ test set. In a within-subject user study (following [13]), 30 participants each rated five samples (from 30 randomly selected tests) on content similarity, instruction alignment, and motion quality using a 5-point Likert scale. Content similarity evaluates whether the edited motion preserves the original meaning of the source motion, while instruction alignment assesses how accurately the edited motion follows the given command. Participants compared our method against four baselines by reviewing randomly shuffled motion outputs. Additionally, we measured performance using data-driven metrics, Frechet Inception Distance (FID), and mean per joint position error (MPJPE), against the labeled motions in the Inter-$\text{MT}^{2}$ test set, following [13]. The results are detailed in §5.3.
@@
We further evaluated our method on three traditional interactive motion tasks: motion-to-text, text-to-motion, and reaction generation, using the combined test sets from InterHuman [22] and Inter-X [39]. Text-motion matching is assessed via top-3 retrieval precision (batch size 32) in the retrieval models’ feature space [29]. Motion quality is measured by the Frechet Inception Distance (FID) and the accuracy of reaction motions is measured by mean per joint position error (MPJPE) in meters. Detailed results are in §5.4.
@@
Since our interactive multi-turn scenarios and tasks, including interactive motion reasoning and motion editing, are novel, there is no exact comparison method. We compare our method against reasonable baselines that handle both motion and texts as input and output.
  • Two-stage approach. We leverage off-the-shelf LLMs and motion-to-text methods. For the motion reasoning task, we convert motions to text via the state-of-the-art motion-to-text model, TM2T [14] and then apply large language models (GPT-4o [27], LLaMA-3.1-8B [9]). For the motion editing task, we first convert the given motions into text descriptions using TM2T, and we concatenate the motion description with editing command texts. We then put the texts to InterGEN [22] to generate modified motions.
  • Extending unified single-human motion model. We adopt a single-human motion-language model, MotionGPT [18], for interactive motions. We consider three variations: (1) $\text{MotionGPT}^{*}$: a modified MotionGPT fine-tuned on interactive motion data; (2) $\text{MotionGPT}^{*}_{I}$: $\text{MotionGPT}^{*}$ enhanced with Inter-$\text{MT}^{2}$ dataset; (3) VIM w/o Inter-$\text{MT}^{2}$: our method fine-tuned with instruction templates, but without Inter-$\text{MT}^{2}$ data.
  • Motion generation baselines. For traditional motion tasks, we benchmarked against interactive motion generation methods, ComMDM [32] and InterGEN [22], and a single human motion generation model, MoMask [15].
@@ 5.2 Motion Reasoning
Methods LLM-Assisted Linguistic Metrics
Coh. $\uparrow$ Align. $\uparrow$ Nat. $\uparrow$ METEOR MAUVE
two-stage approach
TM2T + LLaMA-3.1 3.852 3.050 6.348 0.226 0.009
TM2T + GPT-4o 4.266 3.455 6.790 0.227 0.019
unified approach
$\text{MotionGPT}^{*}$ 1.855 1.303 3.574 0.096 0.005
$\text{MotionGPT}^{*}_{I}$ 3.690 3.160 5.291 0.218 0.417
VIM w/o Inter-$\text{MT}^{2}$ 2.770 2.141 4.968 0.145 0.004
VIM (Ours) 5.252 4.511 6.981 0.260 0.794
@@
In the motion reasoning task, conversations about two interactive motions are examined to assess the model’s ability to deduce past or future events and comprehend the motivations driving the motions. The experimental results in Table 2 demonstrate that our unified model, VIM, outperforms baselines across all LLM-assisted and linguistic metrics. Specifically, VIM achieves improvements with performance increases exceeding 1.9 points in logical coherence, 1.1 points in content alignment, and nearly 0.2 points in naturalness compared to the best two-stage model.
The improved performance of our unified model, VIM, over two-stage approaches, appears to result from two key factors: error accumulation and interpretation ambiguity. Two-stage models can carry over errors if the motion captioning step is inaccurate, undermining content alignment and coherence. In contrast, our unified architecture integrates motion encoding and reasoning in a single framework, minimizing error propagation. Moreover, a single caption may not fully capture multiple interpretations of the same motion, compromising context accuracy in two-stage setups. Our unified approach, however, accounts for these varied interpretations to generate more contextually precise outputs. Figure 5 shows its ability to dynamically adjust interpretations and responses by incorporating context from previous conversations.
@@ 5.3 Motion Editing
Figure 6: User subject study results for motion editing. We plotted the difference only in a post hoc pairwise comparison of the proposed method. * as $0.050.01<p<$ 0.05, ** as $p<0.01$, and *** as $p<0.001$. The error bars represent 95% confidence intervals.
We aimed to validate the hypothesis that users perceive interactive motions edited by our proposed method as more content-consistent, better aligned with instructions, and of higher overall quality. To investigate this, we conducted a user study and analyzed the results using repeated-measures multivariate analysis of variance (MANOVA). The analysis revealed significant effects of the method on user perception across all evaluated dimensions; $F(4)=4.591,p=0.002,\eta^{2}=0.137$ for content similarity, $F(4)=7.134,p=0.000,\eta^{2}=0.197$ for instruction alignment, and $F(4)=4.781,p=0.001,\eta^{2}=0.142$ for motion quality, with all $\alpha=0.05$. The estimated marginal mean of the rated score is reported in Figure 6. The results show that the proposed method had better alignment, quality, and consistency of instruction in other baselines with significant differences.
@@
Table 3: Quantitative results in motion editing task.
@@
During post-hoc pairwise comparisons, VIM significantly outperforms the two-stage model (TM2T [14] with InterGEN [22]) in terms of content similarity ($p=0.017$) and instruction alignment ($p=0.010$). The two-stage model had lower content similarity due to motion-to-text conversion errors causing unintended motions. It also struggled with instruction alignment since InterGEN was trained to generate motions from textual descriptions, limiting its adaptability. In contrast, our unified framework avoids error accumulation and, trained on diverse instructions, demonstrates superior reasoning and adaptability for accurate motion editing and generation. Compared to VIM w/o Inter-$\text{MT}^{2}$, our model significantly improves content similarity ($p=0.010$) and instruction alignment ($p=0.001$), suggesting that excluding Inter-$\text{MT}^{2}$ data hinders motion control. It also outperforms $\text{MotionGPT}^{*}_{I}$ in all metrics, indicating that the baseline’s VQ-based tokenizer struggles to capture precise relative joint positions in two-person motion. Further ablation studies on the motion tokenizer are provided in the supplementary materials. Quantitative evaluations using data-driven metrics, specifically FID and MPJPE (Table 3), further confirmed the superiority of our method over baseline methods, consistent with user study results. Examples of generated edited motions are illustrated in Figure 7.